# The Complexity of Defining and Separating Fixpoint Formulae in Modal Logic

### v1

@@ 4. Unary Case
We first investigate $\mathsf{ML}$-separability over $\mathbb{T}^{1}$, that is, models that are essentially finite or infinite words. Note that satisfiability of $\mu\mathsf{ML}$ over words is PSpace-complete: an upper bound follows, e.g. via the translation to automata and the lower bound is inherited from $\mathsf{LTL}$ [SC85, Theorem 4.1]. This suggests that also definability and separability could be easier. Indeed, we show:
@@
As argued in the previous section, it suffices to construct an $\mathsf{ML}^{n}$-uniform consequence of the NPTA equivalent to $\varphi$, which we do next. {restatable}proppropluniformone Let $\mathcal{A}$ be an NPTA over $\mathbb{T}^{1}$. For every $n\in\mathbb{N}$, an $\mathsf{ML}^{n}$-uniform consequence of $\mathcal{A}$ over $\mathbb{T}^{1}$ can be constructed in time polynomial in $n$ and $|\mathcal{A}|$.
@@ 6. Ternary and Beyond
Intuitively, this means that $\psi_{i},\psi_{i}^{\prime}$ from Lemma 17 enforce in joint $\leftrightarroweq_{\{a\}}^{i}$-consistent models $\mathcal{M},\mathcal{M}^{\prime}$ that $\mathcal{M}$ contains $2^{i}$ points $w_{0},\ldots,w_{2^{i}-1}$ which are all linked to the same point $\widehat{w}$ in $\mathcal{M}^{\prime}$. We exploit this link to synchronize information between the $w_{j}$, following a strategy that has recently been used to show coNExpTime-hardness for interpolant existence in some description logics [AJM+23].
@@
where $\oplus_{n}$ denotes addition modulo $n$. It is well-known that the exponential torus tiling problem, which takes as input $\Delta,H,V,n$ with $n$ in unary and asks whether there is a torus tiling for $\Delta,H,V,2^{n}$, is NExpTime-complete [Für83].
@@
Again, we show the lower bound only for $d=3$; the proof for $d>3$ is analogous. The overall proof strategy is inspired by recently studied interpolant existence problems for description logics [AJM+23], but the details have to be adapted to the present setup. The proof is via a reduction of the word problem for languages recognized by exponentially space bounded, alternating Turing machines, which we recall next.
@@
An ATM $\mathbb{A}$ accepts an input $w$ if there is a computation tree of $\mathbb{A}$ on $w$ in which every branch is infinite. An ATM is called $f$-space bounded for some function $f:\mathbb{N}\to\mathbb{N}$ if in every configuration of every computation tree of $\mathbb{A}$ on an input of length $n$ the non-blank part of the tape is of length at most $f(n)$. It is well-known that there are $2^{n}$-space bounded ATMs which recognize a 2-ExpTime-hard language [CKS81].
@@
The challenge is to enforce that all successor configurations are consistent with $\mathbb{A}$’s transition relation $\Delta$, and we exploit here the bisimilarity of the $w_{i}$ to achieve this. The idea is that, below each $w_{i}$ we coordinate only the $(2^{n}-i)$-th position of each configuration with its successor configuration(s) in the computation tree, as depicted in Figure 2.111The figure contains some additional labeling that will be explained in more detail later. This is possible using auxiliary non-$\sigma$-propositions and due to the fact that we can access the index $i$ of $w_{i}$ via propositions $b_{0},\ldots,b_{n-1}$ mentioned in Lemma 17. Note that this provides sufficient coordination, since, due to the bisimilarity of the $w_{i}$, all the positions will be synchronized. As mentioned above, this overall strategy is similar to one employed in [AJM+23].
@@
For every NPTA $\mathcal{A}$, an NPTA $\mathcal{B}$ recognizing $\mathsf{bisQuot}(\mathcal{L}(\mathcal{A}))$ can be computed in time exponential in the size of $\mathcal{A}$.
@@
Using (18) we prove Proposition 20. It suffices to construct an automaton $\mathcal{B}$ which accepts $\mathcal{M}$ iff $\exists\text{ve}$ wins $\mathcal{G}_{\mathsf{bisQuot}}(\mathcal{M},\mathcal{A})$. By positional determinacy of parity games, to check the latter it suffices to look at positional strategies for $\exists\text{ve}$ in $\mathcal{G}_{\mathsf{bisQuot}}(\mathcal{M},\mathcal{A})$. Using standard techniques we encode $\exists\text{ve}$’s positional strategies for $\mathcal{G}_{\mathsf{bisQuot}}(\mathcal{M},\mathcal{A})$ as colorings of $\mathcal{M}$ with $\mathcal{P}(Q\times Q)$ and construct, in time exponential in $|Q|$, an automaton $\mathcal{B}^{+}$ recognizing models labelled with such winning positional strategies. We then obtain $\mathcal{B}$ recognizing $\mathsf{bisQuot}(\mathcal{L}(\mathcal{A}))$ by projecting out the additional colors $\mathcal{P}(Q\times Q)$ from $\mathcal{B}^{+}$.
@@
We prove the upper bound from Theorem 6.2. Using Theorem 1 compute automata $\mathcal{A},\mathcal{A}^{\prime}$ equivalent to $\varphi,\varphi^{\prime}$. Use Proposition 20 to compute $\mathcal{B},\mathcal{B}^{\prime}$ recognizing $\mathsf{QPL}(\mathcal{L}(\mathcal{A}))$ and $\mathsf{QPL}(\mathcal{L}(\mathcal{A}^{\prime}))$. Recall that any two trees are bisimilar iff they have isomorphic bisimulation quotients. It follows that $\varphi,\varphi^{\prime}$ admit an $\mathsf{ML}^{n}$-separator over $\mathbb{T}^{d}$ iff $\mathcal{A},\mathcal{A}^{\prime}$ are joint $\leftrightarroweq_{\sigma}^{n}$-consistent iff $\mathcal{B},\mathcal{B}^{\prime}$ are joint $\cong^{n}_{\sigma}$ consistent. By Proposition 4, the latter condition holds for all $n\in\mathbb{N}$ iff it holds for $n=|\mathcal{B}|\times|\mathcal{B}^{\prime}|+1$ and this can be tested in time polynomial in $|\mathcal{B}|+|\mathcal{B}^{\prime}|$. Since $\mathcal{A},\mathcal{A}^{\prime}$ are exponential, and $\mathcal{B},\mathcal{B}^{\prime}$ are doubly exponential in the size of $\varphi,\varphi^{\prime}$, this gives the upper bound from Theorem 6.2.
@@ 7. Case Study: Graded Modalities
In this section we apply our techniques and results to the case with graded modal operators. We extend $\mu\mathsf{ML}$ with formulae of the shape $\Diamond_{\geq g}\psi$ and $\Box_{\geq g}\psi$, where $g\in\mathbb{N}$ is a natural number called a grade. We denote with $\mathsf{grML}$ and $\mathsf{gr}\mu\mathsf{ML}$ the extension of $\mathsf{ML}$ and $\mu\mathsf{ML}$, respectively, with such graded modalities. Intuitively, $\Diamond_{\geq g}\psi$ is true in a point $w$ if $w$ has at least $g$ successors satisfying $\psi$ and dually, $\Box_{\geq g}\psi$ is true in $w$ if all but at most $g$ successors satisfy $\psi$ [Fin72, Ott19]. Formally, the semantics of $\mathsf{gr}\mu\mathsf{ML}$ is given by semantic games. These games extend the semantic games used to define the semantics of $\mu\mathsf{ML}$. As in the case with no grades, the game for a formula $\varphi$ and a model $\mathcal{M}$ has $M\times\mathsf{sub}(\varphi)$ as the set of positions, the winning (parity) condition is defined as in the classical case, and so are the moves for all the positions with topmost connective other than the graded modalities. In the classical game, from $(v,\Diamond\theta)$ $\exists\text{ve}$ chooses a child $v^{\prime}$ of $v$ and the next position is $(v^{\prime},\theta)$. In $(v,\Diamond_{\geq k}\theta)$, first $\exists\text{ve}$ chooses a subset $v_{1},...,v_{k}$ of size $k$ of children of $v$, then $\forall\text{dam}$ chooses one of these children $v_{i}$ and the next round starts at $(v_{i},\theta)$. Dually, in $(v,\Box_{k\geq}\theta)$ first $\exists\text{ve}$ picks a subset $v_{1},...,v_{k}$ of at most $k$ $v$’s children, then $\forall\text{dam}$ responds with a choice of some $v^{\prime}$ not in $v_{1},...,v_{k}$ and the next position is $(v^{\prime},\theta)$. Clearly, for any $d\in\mathbb{N}$, the class $\mathbb{T}^{d}$ is $\mathsf{gr}\mu\mathsf{ML}$-definable by the formula $\theta_{d}=\nu x.(\Box x\wedge\Box_{\geq d}\bot)$, which is an additional motivation to study $\mathsf{grML}$ and $\mathsf{gr}\mu\mathsf{ML}$.
@@
We use tools and techniques from the previous sections to solve separability and definability problems in the graded setting. These questions split into two categories, depending on wether grades are allowed in the separating/defining formula or only in the input. Before we inspect these two cases in turn, let us discuss two limitations of this inquiry. First, we only look at the decision problems: deciding existence of a suitable separator/definition. The problem of efficiently constructing one when it exists is a nontrivial task, at least in the case with no grades in the separator, and goes far beyond the scope of this paper. Second, in this section we assume that the grades in formulae are represented in unary. This is admittedly a serious limitation, but is needed for an exponential translation from $\mathsf{gr}\mu\mathsf{ML}$ to automata over $\mathbb{T}^{d}$, generalizing Theorem 1. It is likely that this restriction can be lifted by working with a more succinct automata model suitable for efficiently storing grades, following the approach of [KSV02]. Adapting our proofs to this setting is not immediate, however.
@@ 7.1. With Grades in the Separator
We first investigate $\mathsf{grML}$-definability and $\mathsf{grML}$-separability of $\mathsf{gr}\mu\mathsf{ML}$-formulae. We start by introducing a graded variant of bisimulations called graded bisimulations, which characterize the expressive power of $\mathsf{grML}$ [dR00]. A relation $Z$ between models is a graded bisimulation if it satisfies ($\mathsf{atom}$) and graded variants of the ($\mathsf{back}$) and ($\mathsf{forth}$) conditions of bisimulations. The graded ($\mathsf{forth}$) condition says that if $vZw$ then for every $k\in\mathbb{N}$ and pairwise different children $v_{1},...,v_{k}$ of $v$, there are pairwise different children $w_{1},...,w_{k}$ of $w$ satisfying $v_{i}Zw_{i}$ for all $i\leq k$. The graded ($\mathsf{back}$) condition is symmetric. $Z$ is a $g$-graded bisimulation if the graded ($\mathsf{forth}$) and ($\mathsf{back}$) conditions need to be satisfied only for $k\leq g$. We denote with $\mathcal{M}\leftrightarroweq_{\mathbf{grd}}\mathcal{M}^{\prime}$ (resp., $\mathcal{M}\leftrightarroweq_{\mathbf{grd}(g)}\mathcal{M}^{\prime}$) the fact that there is a graded bisimulation (resp., a $g$-graded bisimulation) between $\mathcal{M}$ and $\mathcal{M}^{\prime}$ that relates their roots. Variants with bounded depth $n$ and/or given signature $\sigma$ are defined and denoted as expected.
@@ 7.2. Without Grades in the Separator
We start with the easier $\mathsf{ML}$-definability of $\mathsf{gr}\mu\mathsf{ML}$-formulae, following a strategy pursued for deciding $\mathsf{ML}$-definability of $\mathsf{grML}$-formulae [KTWZ25]. As an intermediate step, and because it is interesting in its own right, we also look at $\mu\mathsf{ML}$-definability of $\mathsf{gr}\mu\mathsf{ML}$-formulae. It turns out that if a $\mathsf{gr}\mu\mathsf{ML}$-formula is $\mu\mathsf{ML}$-definable then this is already witnessed by the most naive candidate for a definition. Let flattening of a given $\varphi\in\mathsf{gr}\mu\mathsf{ML}$, denoted $\mathsf{flat}(\varphi)$, be a $\mu\mathsf{ML}$-formula obtained from $\varphi$ by replacing every $\Diamond_{\geq k}$ with $\Diamond$ and every $\Box_{\geq k}$ with $\Box$.
@@
Lemma 21 reduces $\mu\mathsf{ML}$-definability of a given $\mathsf{gr}\mu\mathsf{ML}$-formula $\varphi$ to checking the validity of $\varphi\iff\mathsf{flat}(\varphi)$. Moreover, $\varphi$ is $\mathsf{ML}$-definable iff (i) it is equivalent to its flattening $\mathsf{flat}(\varphi)$ and (ii) $\mathsf{flat}(\varphi)$ is $\mathsf{ML}$-definable. Since the validity of $\varphi\iff\mathsf{flat}(\varphi)$ can be decided in exponential time [CGL99, KSV02] and the latter (ii) can be decided via Theorem 3.1, we obtain the following complexities.
@@
Towards the upper bound, assume $\varphi,\varphi^{\prime}\in\mathsf{gr}\mu\mathsf{ML}$. As usual, we only need to look for separators among formulae over $\sigma=\text{sig}(\varphi)\cup\text{sig}(\varphi^{\prime})$. By an immediate generalization of the Equivalence ($\mathsf{Base}$), $\varphi,\varphi^{\prime}$ are not $\mathsf{ML}_{\sigma}^{n}$-separable iff they are joint $\leftrightarroweq_{\sigma}^{n}$-consistent. A standard game-theoretic argument similar to the one in the proof of $3\Rightarrow 4$ in Lemma 7.1 shows that $\varphi,\varphi$ are joint $\leftrightarroweq_{\sigma}^{n}$-consistent over all models iff they are joint $\leftrightarroweq_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$, where $d=g\times(|\varphi|+|\varphi^{\prime}|)$ and $g$ is the greatest grade occurring in $\varphi,\varphi^{\prime}$. Summing up, $\varphi,\varphi^{\prime}$ are not $\mathsf{ML}$-separable iff they are joint $\leftrightarroweq_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$ for all $n\in\mathbb{N}$. To decide the latter we proceed as in the upper bound proof of Theorem 6.2. We first construct NPTA $\mathcal{A},\mathcal{A}^{\prime}$ equivalent to $\varphi,\varphi^{\prime}$ over $d$-ary trees via (an analogue for $\mathsf{gr}\mu\mathsf{ML}$ of) Theorem 1. Then, using Proposition 20 we construct $\mathcal{B}$ recognizing $\mathsf{QPL}(\mathcal{L}(\mathcal{A}))$ and $\mathcal{B}^{\prime}$ recognizing $\mathsf{QPL}(\mathcal{L}(\mathcal{A}^{\prime}))$ and apply Proposition 4 to decide if $\mathcal{B},\mathcal{B}^{\prime}$ are joint $\cong_{\sigma}^{n}$-consistent for all $n$.

### v2

@@ 4. Unary Case
We first investigate $\mathsf{ML}$-separability over $\mathbb{T}^{1}$, that is, models that are essentially finite or infinite words. Note that satisfiability of $\mu\mathsf{ML}$ over words is PSpace-complete: an upper bound follows, e.g. via the translation to automata and the lower bound is inherited from $\mathsf{LTL}$ [36, Theorem 4.1]. This suggests that also definability and separability could be easier. Indeed, we show:
@@
As argued in the previous section, it suffices to construct an $\mathsf{ML}^{n}$-uniform consequence of the NPTA equivalent to $\varphi$, which we do next. {restatable}proppropluniformone Let $\mathcal{A}$ be an NPTA over $\mathbb{T}^{1}$ with alphabet $\Sigma$. For every $n\in\mathbb{N}$, an $\mathsf{ML}^{n}$-uniform consequence of $\mathcal{A}$ over $\mathbb{T}^{1}$ can be constructed in time polynomial in $n$, $|\Sigma|$, and $|\mathcal{A}|$.
@@ 6. Ternary and Beyond
Intuitively, this means that $\psi_{i},\psi_{i}^{\prime}$ from Lemma 17 enforce in joint $\leftrightarroweq_{\{a\}}^{i}$-consistent models $\mathcal{M},\mathcal{M}^{\prime}$ that $\mathcal{M}$ contains $2^{i}$ points $w_{0},\ldots,w_{2^{i}-1}$ which are all linked to the same point $\widehat{w}$ in $\mathcal{M}^{\prime}$. We exploit this link to synchronize information between the $w_{j}$, following a strategy that has recently been used to show coNExpTime-hardness for interpolant existence in some description logics [1].
@@
where $\oplus_{n}$ denotes addition modulo $n$. It is well-known that the exponential torus tiling problem, which takes as input $\Delta,H,V,n$ with $n$ in unary and asks whether there is a torus tiling for $\Delta,H,V,2^{n}$, is NExpTime-complete [17].
@@
Again, we show the lower bound only for $d=3$; the proof for $d>3$ is analogous. The overall proof strategy is inspired by recently studied interpolant existence problems for description logics [1], but the details have to be adapted to the present setup. The proof is via a reduction of the word problem for languages recognized by exponentially space bounded, alternating Turing machines, which we recall next.
@@
An ATM $\mathbb{A}$ accepts an input $w$ if there is a computation tree of $\mathbb{A}$ on $w$ in which every branch is infinite. An ATM is called $f$-space bounded for some function $f:\mathbb{N}\to\mathbb{N}$ if in every configuration of every computation tree of $\mathbb{A}$ on an input of length $n$ the non-blank part of the tape is of length at most $f(n)$. It is well-known that there are $2^{n}$-space bounded ATMs which recognize a 2-ExpTime-hard language [10].
@@
The challenge is to enforce that all successor configurations are consistent with $\mathbb{A}$’s transition relation $\Delta$, and we exploit here the bisimilarity of the $w_{i}$ to achieve this. The idea is that, below each $w_{i}$ we coordinate only the $(2^{n}-i)$-th position of each configuration with its successor configuration(s) in the computation tree, as depicted in Figure 2.111The figure contains some additional labeling that will be explained in more detail later. This is possible using auxiliary non-$\sigma$-propositions and due to the fact that we can access the index $i$ of $w_{i}$ via propositions $b_{0},\ldots,b_{n-1}$ mentioned in Lemma 17. Note that this provides sufficient coordination, since, due to the bisimilarity of the $w_{i}$, all the positions will be synchronized. As mentioned above, this overall strategy is similar to one employed in [1].
@@
For every NPTA $\mathcal{A}$, an NPTA $\mathcal{B}$ in trash normal form recognizing $\mathsf{bisQuot}(\mathcal{L}(\mathcal{A}))$ can be computed in time exponential in the size of $\mathcal{A}$.
@@
Using (18) we prove Proposition 20. It suffices to construct an automaton $\mathcal{B}$ which accepts $\mathcal{M}$ iff $\exists\text{ve}$ wins $\mathcal{G}_{\mathsf{bisQuot}}(\mathcal{M},\mathcal{A})$. By positional determinacy of parity games, to check the latter it suffices to look at positional strategies for $\exists\text{ve}$ in $\mathcal{G}_{\mathsf{bisQuot}}(\mathcal{M},\mathcal{A})$. Using standard techniques we encode $\exists\text{ve}$’s positional strategies for $\mathcal{G}_{\mathsf{bisQuot}}(\mathcal{M},\mathcal{A})$ as colorings of $\mathcal{M}$ with $\mathcal{P}(Q\times Q)$ and construct, in time exponential in $|Q|$, an automaton $\mathcal{B}^{+}$ recognizing models labelled with such winning positional strategies. We then obtain $\mathcal{B}$ recognizing $\mathsf{bisQuot}(\mathcal{L}(\mathcal{A}))$ by projecting out the additional colors $\mathcal{P}(Q\times Q)$ from $\mathcal{B}^{+}$. Finally, we bring $\mathcal{B}$ to trash normal form. To do it in single exponential time we use the following observation, similar to the argument from [13] used to obtain the trash normal form in Theorem 1. Although the number of states of $\mathcal{B}$ is exponential in $|\mathcal{A}|$, the construction can be carried out in such a way that the number of ranks is only polynomial in $|\mathcal{A}|$. This allows to check emptiness of each state of $\mathcal{B}$ in time exponential in $|\mathcal{A}|$.
@@
We prove the upper bound from Theorem 6.2. Using Theorem 1 compute automata $\mathcal{A},\mathcal{A}^{\prime}$ equivalent to $\varphi,\varphi^{\prime}$. Use Proposition 20 to compute automata $\mathcal{B},\mathcal{B}^{\prime}$ in trash normal form which recognize $\mathsf{QPL}(\mathcal{L}(\mathcal{A}))$ and $\mathsf{QPL}(\mathcal{L}(\mathcal{A}^{\prime}))$. Recall that any two trees are bisimilar iff they have isomorphic bisimulation quotients. It follows that $\varphi,\varphi^{\prime}$ admit an $\mathsf{ML}^{n}$-separator over $\mathbb{T}^{d}$ iff $\mathcal{A},\mathcal{A}^{\prime}$ are joint $\leftrightarroweq_{\sigma}^{n}$-consistent iff $\mathcal{B},\mathcal{B}^{\prime}$ are joint $\cong^{n}_{\sigma}$ consistent. By Proposition 4, the latter condition holds for all $n\in\mathbb{N}$ iff it holds for $n=|\mathcal{B}|\times|\mathcal{B}^{\prime}|+1$ and this can be tested in time polynomial in $(|\Sigma|+|\Sigma^{\prime}|)\times(|\mathcal{B}|+|\mathcal{B}^{\prime}|)^{d}$ where $\Sigma=\mathcal{P}(\text{sig}(\varphi))$ is the alphabet of $\mathcal{A}$ and of $\mathcal{B}$, and $\Sigma^{\prime}=\mathcal{P}(\text{sig}(\varphi^{\prime}))$ is the alphabet of $\mathcal{A}^{\prime}$ and of $\mathcal{B}^{\prime}$. Since $\mathcal{A},\mathcal{A}^{\prime}$ are exponential, and $\mathcal{B},\mathcal{B}^{\prime}$ are doubly exponential in the size of $\varphi,\varphi^{\prime}$, this gives the upper bound from Theorem 6.2.
@@ 7. Case Study: Graded Modalities
In this section we apply our techniques and results to the case with graded modal operators. We extend $\mu\mathsf{ML}$ with formulae of the shape $\Diamond_{\geq g}\psi$ and $\Box_{\geq g}\psi$, where $g\in\mathbb{N}$ is a natural number called a grade. We denote with $\mathsf{grML}$ and $\mathsf{gr}\mu\mathsf{ML}$ the extension of $\mathsf{ML}$ and $\mu\mathsf{ML}$, respectively, with such graded modalities. Intuitively, $\Diamond_{\geq g}\psi$ is true in a point $w$ if $w$ has at least $g$ successors satisfying $\psi$ and dually, $\Box_{\geq g}\psi$ is true in $w$ if all but at most $g$ successors satisfy $\psi$ [14, 31]. Formally, the semantics of $\mathsf{gr}\mu\mathsf{ML}$ is given by semantic games. These games extend the semantic games used to define the semantics of $\mu\mathsf{ML}$. As in the case with no grades, the game for a formula $\varphi$ and a model $\mathcal{M}$ has $M\times\mathsf{sub}(\varphi)$ as the set of positions, the winning (parity) condition is defined as in the classical case, and so are the moves for all the positions with topmost connective other than the graded modalities. In the classical game, from $(v,\Diamond\theta)$ $\exists\text{ve}$ chooses a child $v^{\prime}$ of $v$ and the next position is $(v^{\prime},\theta)$. In $(v,\Diamond_{\geq k}\theta)$, first $\exists\text{ve}$ chooses a subset $v_{1},...,v_{k}$ of size $k$ of children of $v$, then $\forall\text{dam}$ chooses one of these children $v_{i}$ and the next round starts at $(v_{i},\theta)$. Dually, in $(v,\Box_{k\geq}\theta)$ first $\exists\text{ve}$ picks a subset $v_{1},...,v_{k}$ of at most $k$ $v$’s children, then $\forall\text{dam}$ responds with a choice of some $v^{\prime}$ not in $v_{1},...,v_{k}$ and the next position is $(v^{\prime},\theta)$. Clearly, for any $d\in\mathbb{N}$, the class $\mathbb{T}^{d}$ is $\mathsf{gr}\mu\mathsf{ML}$-definable by the formula $\theta_{d}=\nu x.(\Box x\wedge\Box_{\geq d}\bot)$, which is an additional motivation to study $\mathsf{grML}$ and $\mathsf{gr}\mu\mathsf{ML}$.
@@
We use tools and techniques from the previous sections to solve separability and definability problems in the graded setting. These questions split into two categories, depending on wether grades are allowed in the separating/defining formula or only in the input. Before we inspect these two cases in turn, let us discuss two limitations of this inquiry. First, we only look at the decision problems: deciding existence of a suitable separator/definition. The problem of efficiently constructing one when it exists is a nontrivial task, at least in the case with no grades in the separator, and goes far beyond the scope of this paper. Second, in this section we assume that the grades in formulae are represented in unary. This is admittedly a serious limitation, but is needed for an exponential translation from $\mathsf{gr}\mu\mathsf{ML}$ to automata over $\mathbb{T}^{d}$, generalizing Theorem 1. It is likely that this restriction can be lifted by working with a more succinct automata model suitable for efficiently storing grades, following the approach of [27]. Adapting our proofs to this setting is not immediate, however.
@@ 7.1. With Grades in the Separator
We first investigate $\mathsf{grML}$-definability and $\mathsf{grML}$-separability of $\mathsf{gr}\mu\mathsf{ML}$-formulae. We start by introducing a graded variant of bisimulations called graded bisimulations, which characterize the expressive power of $\mathsf{grML}$ [12]. A relation $Z$ between models is a graded bisimulation if it satisfies ($\mathsf{atom}$) and graded variants of the ($\mathsf{back}$) and ($\mathsf{forth}$) conditions of bisimulations. The graded ($\mathsf{forth}$) condition says that if $vZw$ then for every $k\in\mathbb{N}$ and pairwise different children $v_{1},...,v_{k}$ of $v$, there are pairwise different children $w_{1},...,w_{k}$ of $w$ satisfying $v_{i}Zw_{i}$ for all $i\leq k$. The graded ($\mathsf{back}$) condition is symmetric. $Z$ is a $g$-graded bisimulation if the graded ($\mathsf{forth}$) and ($\mathsf{back}$) conditions need to be satisfied only for $k\leq g$. We denote with $\mathcal{M}\leftrightarroweq_{\mathbf{grd}}\mathcal{M}^{\prime}$ (resp., $\mathcal{M}\leftrightarroweq_{\mathbf{grd}(g)}\mathcal{M}^{\prime}$) the fact that there is a graded bisimulation (resp., a $g$-graded bisimulation) between $\mathcal{M}$ and $\mathcal{M}^{\prime}$ that relates their roots. Variants with bounded depth $n$ and/or given signature $\sigma$ are defined and denoted as expected.
@@ 7.2. Without Grades in the Separator
We start with the easier $\mathsf{ML}$-definability of $\mathsf{gr}\mu\mathsf{ML}$-formulae, following a strategy pursued for deciding $\mathsf{ML}$-definability of $\mathsf{grML}$-formulae [26]. As an intermediate step, and because it is interesting in its own right, we also look at $\mu\mathsf{ML}$-definability of $\mathsf{gr}\mu\mathsf{ML}$-formulae. It turns out that if a $\mathsf{gr}\mu\mathsf{ML}$-formula is $\mu\mathsf{ML}$-definable then this is already witnessed by the most naive candidate for a definition. Let flattening of a given $\varphi\in\mathsf{gr}\mu\mathsf{ML}$, denoted $\mathsf{flat}(\varphi)$, be a $\mu\mathsf{ML}$-formula obtained from $\varphi$ by replacing every $\Diamond_{\geq k}$ with $\Diamond$ and every $\Box_{\geq k}$ with $\Box$.
@@
Lemma 21 reduces $\mu\mathsf{ML}$-definability of a given $\mathsf{gr}\mu\mathsf{ML}$-formula $\varphi$ to checking the validity of $\varphi\iff\mathsf{flat}(\varphi)$. Moreover, $\varphi$ is $\mathsf{ML}$-definable iff (i) it is equivalent to its flattening $\mathsf{flat}(\varphi)$ and (ii) $\mathsf{flat}(\varphi)$ is $\mathsf{ML}$-definable. Since the validity of $\varphi\iff\mathsf{flat}(\varphi)$ can be decided in exponential time [9, 27] and the latter (ii) can be decided via Theorem 3.1, we obtain the following complexities.
@@
Towards the upper bound, assume $\varphi,\varphi^{\prime}\in\mathsf{gr}\mu\mathsf{ML}$. As usual, we only need to look for separators among formulae over $\sigma=\text{sig}(\varphi)\cup\text{sig}(\varphi^{\prime})$. By an immediate generalization of the Equivalence ($\mathsf{Base}$), $\varphi,\varphi^{\prime}$ are not $\mathsf{ML}_{\sigma}^{n}$-separable iff they are joint $\leftrightarroweq_{\sigma}^{n}$-consistent. A standard game-theoretic argument similar to the one in the proof of $3\Rightarrow 4$ in Lemma 7.1 shows that $\varphi,\varphi$ are joint $\leftrightarroweq_{\sigma}^{n}$-consistent over all models iff they are joint $\leftrightarroweq_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$, where $d=g\times(|\varphi|+|\varphi^{\prime}|)$ and $g$ is the greatest grade occurring in $\varphi,\varphi^{\prime}$. Summing up, $\varphi,\varphi^{\prime}$ are not $\mathsf{ML}$-separable iff they are joint $\leftrightarroweq_{\sigma}^{n}$-consistent over $\mathbb{T}^{d}$ for all $n\in\mathbb{N}$. To decide the latter we proceed as in the upper bound proof of Theorem 6.2. We first construct NPTA $\mathcal{A},\mathcal{A}^{\prime}$ equivalent to $\varphi,\varphi^{\prime}$ over $d$-ary trees via (an analogue for $\mathsf{gr}\mu\mathsf{ML}$ of) Theorem 1. Then, using Proposition 20 we construct $\mathcal{B}$ recognizing $\mathsf{QPL}(\mathcal{L}(\mathcal{A}))$ and $\mathcal{B}^{\prime}$ recognizing $\mathsf{QPL}(\mathcal{L}(\mathcal{A}^{\prime}))$ and apply Proposition 4 to decide whether $\mathcal{B},\mathcal{B}^{\prime}$ are joint $\cong_{\sigma}^{n}$-consistent for all $n$.